OpenStudy (anonymous):

3/x−1+8/x=3( base on a quadratic equation) Ex: 2^2-5x+2=0 (2x-1)(x-2)=0 x+1/x= 5/2 2x-1=0 or x-2=0 the only problem: I get then quadratic equation) -X^2 -11x-10

OpenStudy (mertsj):

\[\frac{3}{x-1}+\frac{8}{x}=3\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

when I start working though I got little confuse of this eqatuion:\[ -X^2+1x-10\]

OpenStudy (anonymous):

I mean this X^2+11x−10

OpenStudy (anonymous):

I mean agian -X^2+11x-10. sorry

OpenStudy (mertsj):